Lithium metal reacts with water to produce aqueous lithium hydroxide and hydrogen gas. The balanced equation for this reaction is shown below. 2Li(s) + 2H 2 O(l) → 2LiOH(aq) + H 2 (g) a) Calculate the mass of H 2 (molar mass 2.016 g mol –1 ) gas that can form in the reaction of 80.6 g of Li (molar mass 6.940 g mol –1 ) with excess H 2 O.
Answer #1

2Li + 2H2O = 2LiOH + H2

From the above reaction it is clear that 2 moles of Li will produce 1 mole of H2.

80.6 g of Li = 80.6/ 6.940 = 11.6 mole [ molar mass= 6.940 g.mol-1]

So, 11.6 moles of Li will produce 11.6/2 =5.8 moles of H2

1 mole of H2 = 2.016 g H2

5.8 moles of H2 = (5.82.016) = 11.7 g of H2

so, 11.7 g of H2 will produce.